### v4

@@ 2.2 Strict Saddle Points
The Center Stable Manifold (CSM) Theorem plays a crucial role in deriving the avoidance of strict saddle points. The CSM theorem establishes a favorable property for the fixed-point iteration of a $C^{1}$ local diffeomorphism map: the set of initial points around a fixed point, which converge to the fixed point with at least one eigenvalue of $DT$ larger than $1$ in magnitude, has Lebesgue measure zero.
Theorem 7.
@@
(The Center Stable Manifold Theorem) Let 00 be a fixed point for the $C^{1}$ local diffeomorphism $T:U\rightarrow\mathbb{R}^{n}$ where $U$ is a neighborhood of zero in $\mathbb{R}^{n}$. Let $E^{s}\oplus E^{c}\oplus E^{u}$ be the invariant splitting of $\mathbb{R}^{n}$ into the generalized eigenspaces of $DT(0)$ corresponding to eigenvalues of absolute value less than one, equal to one, and greater than one. There exists a $C^{1}$ embedded disk $W_{loc}^{cs}$ that is tangent to $E_{s}\oplus E_{c}$ at 0 and a neighborhood $B$ around 0 such that $T(W_{loc}^{cs})\cap B\subset W_{loc}^{cs}$. In addition, if $T^{k}(x)\in B$ for all $k\geq 0$, then $x\in W_{loc}^{cs}$.
@@
If $\nabla T(0)$ is invertible, then $T$ is locally a diffeomorphism around 0. The CSM theorem only guarantees the avoidance of strict saddle point in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points that converge to any unstable fixed point has measure zero. This requires that the map has the so-called 0-property, meaning the preimage of a measure zero set under this map still has measure zero. This result was established in [47, Theorem 2] for a global diffeomorphism map and then for a global lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous) in [23]. Moreover, to extend this property from a particular stationary point $x^{*}$ to the set of all stationary points, this can be shown using the fact that $\mathbb{R}^{n}$ is second countable no matter if the number of stationary points is countable (see [23, Corollary 2.12] for a standard proof). We list their results below.
Theorem 8.
@@
[47, Theorem 2] Let $g$ be a $C^{1}$ mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $det(DT(x))\neq 0$ for all $x\in\mathbb{R}^{n}$. Let $\mathcal{U}_{T}$ consist of all unstable fixed points $x$ of $T$ Then the set of initial points that converge to an unstable fixed point has measure zero, $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
Corollary 1.
@@
[23, Corollary 12] Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a lipeomorphism and let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ at which the Jacobian $\nabla T(x)$ is invertible. Then, the set of initial points attracted by such fixed points $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim\limits_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
@@
    Interpret the iterative algorithm as a fixed point iteration, and build the equivalence of the unstable fixed points and the strict saddle points.
@@
    Show that the fixed point iteration is locally a $C^{1}$ local diffeomorphism around a strict saddle point (e.g., a $C^{1}$ smooth map with invertible Jacobian at the fixed point), so that the CSM theorem guarantees the avoidance of converging to the strict saddle point.
@@ 3 Damped Iterative Reweighted $\ell_{1}$ Algorithms
$x^{k+1}=S(x^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i})|y_{i}|,$
where $w(x_{i})=r^{\prime}(|x_{i}|)$.
This algorithm may quickly converge to local minimizers due to the rapid growth of the weight values, especially for the case where $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$ can be quickly encountered. To address this issue, a common approach is to add a perturbation $\epsilon_{i}>0$ to each $r$ and drive $\epsilon_{i}$ to 0 during the iteration, resulting in the following approximation $F(x,\epsilon)$ of $F(x)$
$\min\limits_{x,\epsilon\in\mathbb{R}^{n}}F(x,\epsilon):=f(x)+\lambda\sum\limits_{i=1}^{n}r(|x_{i}|+\epsilon_{i}).$ (17)
Here we also treat $\epsilon$ as variables, so that the algorithm updates $x$ and $\epsilon$ by $(x^{k+1},\epsilon^{k+1})=S(x^{k},\epsilon^{k})$
$\begin{cases}x^{k+1}=S^{x}(x^{k},\epsilon^{k})=\arg\min\limits_{y}G(y;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum\limits_{i=1}^{n}w_{i}^{k}|y_{i}|,\\ \epsilon^{k+1}=S^{\epsilon}(x^{k},\epsilon^{k})=\mu\epsilon^{k}\end{cases}$
with $w_{i}^{k}=w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$ and $\mu\in(0,1)$.
The subproblem solution can be given explicitly using the solft-thresholding operator,
$S^{x}(x^{k},\epsilon)=\mathbb{S}_{\lambda w^{k}/\beta}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})),$ (18)
@@
$S^{x}_{i}(x^{k},\epsilon^{k})=\begin{cases}x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ 0,&\text{otherwise }.\end{cases}$ (19)
To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a stationary point and a Lipeomorphism globally. However, this is not true in our case since $S$ is generally noninvertible. In particular, $S$ is not a bijection, as it maps multiple values to 00. Motivated by the technique used in [23], we propose to use the damped operator: $T=(1-\alpha)I+\alpha S$, which can be written explicitly as
$(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S^{x}(x^{k},\epsilon^{k})\\ (1-\alpha(1-\mu))\epsilon^{k}\end{bmatrix}$ (20)
@@
Proof.
Suppose $x^{*}$ is a stationary point. It follows from (6) that if $x_{i}^{*}>0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)]=x_{i}^{*}>0$, implying $S_{i}^{x}(x^{*},0)=x_{i}^{*}$. This result also holds true by the same argument if $x_{i}^{*}<0$.
As for $x_{i}^{*}=0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)]<0$ by (7). Same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$. Therefore, $S_{i}^{x}(x^{*},0)=0=x_{i}^{*}$. In general, we have shown that $x^{*}=S^{x}(x^{*},0)$ in all cases.
Conversely, suppose $x^{*}=S^{x}(x^{*},0)$. If $x_{i}^{*}>0$, it must be true that $x_{i}^{*}=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]>0$, implying $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]$. Same argument yields $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|x_{i}^{*}|)]$ if $x_{i}^{*}<0$. Therefore, $\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)=0$ for $i\in\mathcal{I}^{*}$.
@@
The rest part of the statement of this proposition is obviously true by noticing that $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$ implies that $\epsilon^{*}=0$. ∎
1:Input $\beta>\alpha L_{f}/2$, $\alpha\in(0,1)$, $x_{0}$, and $\epsilon_{0}>0$.
@@ 3.1 Convergence analysis
Given the initial point $(x^{0}$, $\epsilon^{0})$ and $F^{0}=F(x^{0},\epsilon^{0})$, the level set $\mathcal{L}(F^{0}):=\{(x,\epsilon)\mid F(x,\epsilon)\leq F^{0}\}$ is bounded.
This assumption means $f$ is Lipschitz continuous and Lipschitz differentiable on $\mathcal{L}(F^{0})$ with constant $L_{f}$, i.e., $\forall x,y\in\mathbb{R}^{n}$ (we use the same constant for $f$ and $\nabla f$ for simplicity).
$|f(x)-f(y)|\leq L_{f}\|x-y\|\text{ \ and \ }\|\nabla f(x)-\nabla f(y)\|\leq L_{f}\|x-y\|.$ (21)
This assumption also implies that if $r^{\prime}(|0^{+}|)=\infty$, then for any stationary point $x^{*}$, $|x_{i}^{*}|\geq(r^{\prime})^{-1}(L_{f}/\lambda),\forall i\in\mathcal{I}^{*}$ by (6). Therefore, we know $r^{\prime\prime}(|x_{i}^{*}|)>0$ is bounded above and this is also the case for $r^{\prime}(|0^{+}|)<\infty$. Overall, we can conclude that for any stationary point $x^{*}\in\mathcal{L}(F^{0})$, there exists $\rho$ such that
$\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho.$ (22)
We first show the monotonicity of $F(x,\epsilon)$. Define the reductions of $G$ and $F$ caused by $x^{k+1}$ as
$\Delta F(x^{k+1},\epsilon^{k+1})=$ $\ F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k+1}),$
@@
Proposition 10.
Suppose Assumptions 1-3 hold. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 10 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{f}}{2}$. It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and that the reduction satisfies
$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
Moreover, $\lim\limits_{k\rightarrow\infty}\|x^{k+1}-x^{k}\|=0$ and there exists $C>0$ such that
$\|x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})\|_{\infty}\leq C,$ (23)
for any $k\in\mathbb{N}$.
Proof.
@@
$r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})\leq r(|x_{i}^{k}|+\epsilon_{i}^{k})+r^{\prime}(|x_{i}^{k}|+\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
@@
$\sum\limits_{i=1}^{n}r(|x_{i}^{k+1}|+\epsilon_{i}^{k})-\sum\limits_{i=1}^{n}r(|x_{i}^{k}|+\epsilon_{i}^{k})\leq\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$ (24)
@@
$\Delta F(x^{k+1},\epsilon^{k+1})\geq F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k})$ (26)
$\geq$ $\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(|x_{i}^{k}|-|x_{i}^{k+1}|)$
$=$ $\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{f}}{2}\|x^{k}-x^{k+1}\|^{2},$
@@
For each subproblem, $y^{k}$ satisfies the optimal condition
$\nabla f(x^{k})+\lambda W(x^{k},\epsilon^{k})\xi^{k}+\beta(y^{k}-x^{k})=0.$ (27)
with $\xi^{k}\in\partial\|y^{k}\|_{1}$ and $W(x^{k},\epsilon^{k}):={\rm diag}(w(x_{1}^{k},\epsilon_{1}^{k}),...,w(x_{n}^{k},\epsilon_{n}^{k}))$. Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimal condition (27), the following holds
$\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$ (28)
@@
where the first inequality is by $\alpha\in(0,1)$ and the second is by the convexity of ||$\cdot$|| ⋅ |. Combine (26), (37) and get
$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
@@
$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
@@
$(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (23) holds for any $k\in\mathbb{N}$. ∎

### v5

@@ 2.3 Center Stable Manifold Theorem
The Center Stable Manifold Theorem (CSM) plays a crucial role in the avoidance of strict saddle points. The CSM theorem establishes a favorable property for the fixed-point iteration of a $C^{1}$ local diffeomorphism map: the set of initial points around a fixed point, from which an algorithm converges to the fixed point with at least one eigenvalue of $DT$ greater than $1$ in magnitude, has Lebesgue measure zero.
Theorem 2.13.
@@
(The Center Stable Manifold Theorem) Let 00 be a fixed point for the $C^{1}$ local diffeomorphism $T:U\rightarrow\mathbb{R}^{n}$ where $U$ is a neighborhood of zero in $\mathbb{R}^{n}$. Let $E^{s}\oplus E^{c}\oplus E^{u}$ be the invariant split of $\mathbb{R}^{n}$ into the generalized eigenspaces of $DT(0)$ corresponding to eigenvalues of absolute value less than one, equal to one and greater than one. There exists a $C^{1}$ embedded disk $W_{loc}^{cs}$ that is tangent to $E_{s}\oplus E_{c}$ at 0 and a neighborhood $B$ around 0 such that $T(W_{loc}^{cs})\cap B\subset W_{loc}^{cs}$. In addition, if $T^{k}(x)\in B$ for all $k\geq 0$, then $x\in W_{loc}^{cs}$.
@@
If $DT(0)$ is invertible, then $T$ is locally a diffeomorphism around 0. The CSM theorem only guarantees the avoidance of strict saddle points in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points from which an algorithm converges to any unstable fixed point has a measure zero. This requires that the map has the so-called 0-property, meaning the preimage of a measure zero set under this map still has measure zero. This result was established in [24, Theorem 2] for a global diffeomorphism map and then for a global lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous) in [15]. Moreover, to extend this property from a particular stationary point $x^{*}$ to the set of all stationary points, this can be shown using the fact that $\mathbb{R}^{n}$ is second countable no matter if the number of stationary points is countable (see [15, Corollary 2.12] for a standard proof). We list their results below.
Theorem 2.14.
@@
[24, Theorem 2] Let $g$ be a $C^{1}$ mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $det(DT(x))\neq 0$ for all $x\in\mathbb{R}^{n}$. Let $\mathcal{U}_{T}$ consist of all unstable fixed points $x$ of $T$ Then the set of initial points from which an algorithm converges to an unstable fixed point has measure zero, $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
Corollary 2.15.
@@
[15, Corollary 12] Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a lipeomorphism and let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ at which the Jacobian $DT(x)$ is invertible. Then, the set of initial points attracted by such fixed points $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim\limits_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
@@
    Interpret the iterative algorithm as a fixed point iteration and build the equivalence of the unstable fixed points and the strict saddle points.
@@
    Show that the fixed point iteration is locally a $C^{1}$ local diffeomorphism around a strict saddle point (e.g., a $C^{1}$ smooth map with invertible Jacobian at the fixed point) so that the CSM theorem guarantees the avoidance of converging to a strict saddle point.
@@ 3 Damped Iterative Reweighted $\ell_{1}$ Algorithms
$x^{k+1}=S^{x}(x^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}(y-x^{k})+\tfrac{\beta}{2}\|y-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}\tilde{w}^{k}_{i}|y_{i}|,$
where weight $\tilde{w}_{i}^{k}=r^{\prime}(|x_{i}^{k}|)$.
This algorithm may quickly converge to local minima due to the rapid growth of the weight values, especially for the case where $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$. To address this issue, a common approach is to add a perturbation $\epsilon_{i}>0$ to each $r$ and drive $\epsilon_{i}$ to 0 during the iteration, resulting in the following approximation $F(x,\epsilon)$ of $F(x)$
(16) $\min\limits_{x\in\mathbb{R}^{n},\epsilon\in\mathbb{R}^{n}_{+}}F(x,\epsilon):=f(x)+\lambda\sum\limits_{i=1}^{n}r(|x_{i}|+\epsilon_{i}).$
Here we treat $\epsilon$ also as variables, so that the algorithm updates $x$ and $\epsilon$ by
$(x^{k+1};\epsilon^{k+1})=S(x^{k},\epsilon^{k}),$
where $S=[S^{x};S^{\epsilon}]$ and
$x^{k+1}$ $=S^{x}(x^{k},\epsilon^{k})=\arg\min\limits_{y}G(y;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum\limits_{i=1}^{n}w_{i}^{k}|y_{i}|,$
$\epsilon^{k+1}$ $=S^{\epsilon}(x^{k},\epsilon^{k})=\mu\epsilon^{k}$
with $w_{i}^{k}:=w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$ and $\mu\in(0,1)$.
The subproblem solution can be given using the soft-thresholding operator,
(17) $S^{x}(x^{k},\epsilon)=\mathbb{S}_{\lambda w^{k}/\beta}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})),$
@@
(18) $S^{x}_{i}(x^{k},\epsilon^{k})=\begin{cases}x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ 0,&\text{otherwise }.\end{cases}$
To take advantage of the CSM theorem, the map $S$ needs to be a local diffeomorphism at a stationary point and a lipeomorphism globally. However, this is not true in our case, since $S$ is generally noninvertible. In particular, $S$ is not a bijection, as it maps multiple values to 00. Motivated by the technique used in [15], we propose to use the damped operator: $T=(1-\alpha)I+\alpha S$, which can be written explicitly as
(19) $(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S^{x}(x^{k},\epsilon^{k})\\ (1-\alpha(1-\mu))\epsilon^{k}\end{bmatrix}$
@@
Proof 3.2.
Suppose $x^{*}$ is a stationary point. It follows from (5) that if $x_{i}^{*}>0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)]=x_{i}^{*}>0$, implying $S_{i}^{x}(x^{*},0)=x_{i}^{*}$. This result also holds by the same argument if $x_{i}^{*}<0$.
As for $x_{i}^{*}=0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)]<0$ by (6). The same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$. Therefore, $S_{i}^{x}(x^{*},0)=0=x_{i}^{*}$. In general, we have shown that $x^{*}=S^{x}(x^{*},0)$ in all cases.
Conversely, suppose $x^{*}=S^{x}(x^{*},0)$. If $x_{i}^{*}>0$, it must be true that $x_{i}^{*}=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]>0$, implying $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]$. The same argument yields $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|x_{i}^{*}|)]$ if $x_{i}^{*}<0$. Therefore, $\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)=0$ for $i\in\mathcal{I}^{*}$.
@@
The rest part of the statement of this proposition is obviously true by noticing that $(x^{*};\epsilon^{*})=T(x^{*},\epsilon^{*})$ implies that $\epsilon^{*}=0$.
1:  Input $\beta>\alpha L_{\nabla f}/2$, $\alpha\in(0,1)$, $x^{0}$, $\epsilon^{0}>0$ and $\mu\in(0,1)$.
@@ 3.1 Convergence analysis
Given the initial point $(x^{0}$, $\epsilon^{0})$ and $F^{0}=F(x^{0},\epsilon^{0})$, the level set $\mathcal{L}(F^{0}):=\{x\mid F(x)\leq F^{0}\}$ is bounded.
This assumption means that there exists $C>0$ such that
(20) $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C,\quad\forall x\in\mathcal{L}(F^{0}).$
It also means $f$ is Lipschitz continuous and differentiable with a Lipschitz gradient on $\mathcal{L}(F^{0})$. Hence, there exist constants $L_{f}$ and $L_{\nabla f}$, such that $\forall x,y\in\mathcal{L}(F^{0})$,
(21) $|f(x)-f(y)|\leq L_{f}\|x-y\|\text{ and }\|\nabla f(x)-\nabla f(y)\|\leq L_{\nabla f}\|x-y\|.$
Moreover, (21) implies that if $r^{\prime}(|0^{+}|)=\infty$, then for any stationary point $x^{*}\in\mathcal{L}(F^{0})$, $|x_{i}^{*}|\geq(r^{\prime})^{-1}(L_{f}/\lambda),\forall i\in\mathcal{I}^{*}$ by (5) and $r^{\prime\prime}(|x_{i}^{*}|)>0$ is bounded above. As for $r^{\prime}(|0^{+}|)<\infty$, it holds naturally that $r^{\prime\prime}(|x_{i}^{*}|)$ is bounded above for $i\in\mathcal{I}^{*}$.
Overall, we can conclude that for any stationary point $x^{*}\in\mathcal{L}(F^{0})$, there exists $\rho>0$ such that
(22) $\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho.$
We first show the monotonicity of $F(x,\epsilon)$. Considering the model $G$ at $(x^{k},\epsilon^{k})$
$G(x;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}x+\frac{\beta}{2}\|x^{k}-x\|^{2}+\lambda\sum_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})|x_{i}|$
with minimum $y^{k}$, its reduction caused by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ is defined as
@@
The reduction of $F$ caused by $(x^{k+1},\epsilon^{k+1})$ with $\epsilon^{k+1}=(1-\alpha(1-\mu))\epsilon^{k}$ is defined as
$\Delta F(x^{k+1},\epsilon^{k+1})=$ $\ F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k+1}).$
Proposition 3.3.
Suppose Assumptions 1-3 hold. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 1 with $\alpha\in(0,1)$ and $\beta\geq\frac{\alpha L_{\nabla f}}{2}$. It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and that the reduction satisfies
$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2},$
which means $\lim\limits_{k\rightarrow\infty}\|x^{k+1}-x^{k}\|=0$. Moreover, $\{x^{k}\}\subset\mathcal{L}(F^{0})$ and (20) holds for any $x=x^{k},k\in\mathbb{N}$.
Proof 3.4.
@@
$r(|x_{i}^{k+1}|+\epsilon_{i}^{k})\leq r(|x_{i}^{k}|+\epsilon_{i}^{k})+r^{\prime}(|x_{i}^{k}|+\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
@@
(23) $\sum\limits_{i=1}^{n}r(|x_{i}^{k+1}|+\epsilon_{i}^{k})-\sum\limits_{i=1}^{n}r(|x_{i}^{k}|+\epsilon_{i}^{k})\leq\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
@@
(25) $\Delta F(x^{k+1},\epsilon^{k+1})\geq F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k})$
$\geq$ $\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{\nabla f}}{2}\|x^{k}-x^{k+1}\|^{2}+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(|x_{i}^{k}|-|x_{i}^{k+1}|)$
$=$ $\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{\nabla f}}{2}\|x^{k}-x^{k+1}\|^{2},$
@@
For each subproblem, $y^{k}$ satisfies the optimality condition
(26) $\nabla f(x^{k})+\lambda W(x^{k},\epsilon^{k})\xi^{k}+\beta(y^{k}-x^{k})=\boldsymbol{0}.$
with $\xi^{k}\in\partial\|y^{k}\|_{1}$ and $W(x^{k},\epsilon^{k}):={\rm diag}(w(x_{1}^{k},\epsilon_{1}^{k}),...,w(x_{n}^{k},\epsilon_{n}^{k}))$. Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimality condition (26), the following holds
(27) $\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$
@@
where the first inequality is by $\alpha\in(0,1)$ and the second is by the convexity of ||$\cdot$|| ⋅ |. Combining (25) and (27), we get
$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
@@
$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
@@
$(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$.
Moreover, $F(x^{k})\leq F(x^{k},\epsilon^{k})\leq F(x^{0},\epsilon^{0})$, meaning $\{x^{k}\}\subset\mathcal{L}(F^{0})$. Therefore, (20) holds for all $x=x^{k},k\in\mathbb{N}$.